Suppose * is an associative and commutative binary operation of a set S. Show that H = { a is an element of S | a*a=a} is closed under the operation *. 
1
Expert's answer
2021-02-12T10:34:21-0500

Let "*" be an associative commutative binary operation on a set "S". Let us show that "H=\{a\in S\ |\ a*a=a\}" is closed under operation "*." Let "a,b\in H". Then "a*a=a" and "b*b=b" . It follows from commutativity and associativity of "*" that"(a*b)*(a*b)= a*(b*(a*b))= a*((b*a)*b)= a*((a*b)*b)= a*(a*(b*b))= a*(a*b)= (a*a)*b=a*b."

Therefore, "a*b\in H", and "H" is closed under operation "*."